# PlantDeBERTa: An Open Source Language Model for Plant Science

### v3

@@ 1 Introduction
Language Models (LMs) [18, 2, 1] have emerged as pivotal tools across diverse NLP tasks, including machine translation, sentiment analysis, and especially named entity recognition (NER) [21, 22]. Their ability to generate and contextualize language representations has transformed how we extract structured information from textual corpora, enabling downstream applications in information retrieval, automated summarization, and scientific discovery [7]. In specialized fields such as medicine and environmental science [15, 23], domain-adapted models have demonstrated the power of NER in extracting actionable insights from unstructured text, supporting applications from drug discovery to ecological monitoring [22, 12].
Traditional NER methods—often rule-based or statistical—fail to generalize well across biological subdomains due to high variability in terminology and contextual usage. Transformer-based models have addressed these limitations by capturing deeper contextual relationships between entities, improving classification and disambiguation performance across various biomedical and scientific domains [9, 8, 3]. Nevertheless, plant science remains underrepresented in both model design and data curation efforts.
@@ 2.3 NLP in Plant Science, Agriculture, and Biotechnology
Significant progress has been made in applying NLP techniques to plant sciences, agriculture, and biotechnology, particularly for NER and knowledge extraction tasks in these domains. One notable application is plant disease and pest management: [19] developed an ALBERT-based NER model for tomato leaf diseases, achieving recall rates above 95% and storing extracted knowledge in a Neo4j graph database. This enabled a digital pest and disease diagnostic tool, demonstrating the practical impact of NLP-powered knowledge graphs in agriculture.
To facilitate such advances, domain-specific corpora and benchmarks are being created for plant science. The Plant Science Knowledge Graph Corpus (PICKLE) is a recent dataset containing 250 annotated plant science abstracts [11]. Using this resource, models trained on PICKLE achieved strong in-domain performance, surpassing even biomedical-trained models. Similarly, the ORKG Agri-NER initiative is defining standard ontologies and entity types for agriculture research text to enable FAIR (Findable, Accessible, Interoperable, Reusable) knowledge extraction [17]. These resources and domain-focused approaches are pushing the frontiers of NLP in plant and agricultural sciences, enhancing entity extraction and knowledge discovery.

### v4

@@ 1 Introduction
Language Models (LMs) [18, 2, Aghaei2022SecureBERT] have emerged as pivotal tools across diverse NLP tasks, including machine translation, sentiment analysis, and especially named entity recognition (NER) [21, 22]. Their ability to generate and contextualize language representations has transformed how we extract structured information from textual corpora, enabling downstream applications in information retrieval, automated summarization, and scientific discovery [7]. In specialized fields such as medicine and environmental science [15, 23], domain-adapted models have demonstrated the power of NER in extracting actionable insights from unstructured text, supporting applications from drug discovery to ecological monitoring [22, 12].
Traditional NER methods—often rule-based or statistical—fail to generalize well across biological subdomains due to high variability in terminology and contextual usage. Transformer-based models have addressed these limitations by capturing deeper contextual relationships between entities, improving classification and disambiguation performance across various biomedical and scientific domains [9, 8, Chandak2022BuildingAK]. Nevertheless, plant science remains underrepresented in both model design and data curation efforts.
@@ 2.3 NLP in Plant Science, Agriculture, and Biotechnology
Significant progress has been made in applying NLP techniques to plant sciences, agriculture, and biotechnology, particularly for NER and knowledge extraction tasks in these domains. One notable application is plant disease and pest management: Wang et al. developed an ALBERT–BiLSTM–CRF NER model for tomato leaf pests and diseases, achieving $\sim$95% recall and storing extracted knowledge in a Neo4j graph database; this enabled a digital diagnostic tool for end users [19].
To facilitate such advances, domain-specific corpora and benchmarks are being created for plant science. The Plant Science Knowledge Graph Corpus (PICKLE) is a recent dataset containing 250 annotated plant science abstracts [Lotreck2024PICKLE]. Using this resource, models trained on PICKLE achieved strong in-domain performance, surpassing even biomedical-trained models. Similarly, the ORKG Agri-NER initiative is defining standard ontologies and entity types for agriculture research text to enable FAIR (Findable, Accessible, Interoperable, Reusable) knowledge extraction [17]. These resources and domain-focused approaches are pushing the frontiers of NLP in plant and agricultural sciences, enhancing entity extraction and knowledge discovery.